OpenStudy (anonymous):

How to state the period, phase shift, and vertical displacement for each equation: y=tan(x+60) y=2cot(2x-45°)

OpenStudy (anonymous):

@esamalaa

OpenStudy (anonymous):

You, ma'am, are answering the first question right.

OpenStudy (anonymous):

For the second one: Period = 180/2 = 90 Phase shift = 45/2 = 22.5 Vertical displacement = 0 Quite simple, really. Where do you get stuck?